Bond Coupon Rate

Cranston Industries just issued $1000 par 30 year bonds. The bonds sold for $1107.20 and pay interest semi annually. Investors require a rate of 7.75% on the bonds.What's the bonds coupon rate?
Answer #1
We have FV = 1000, Sale price = PV = 1107.20,
Semiannual bond. SO each year has 2 periods. SO nper = 30 yr = 30*2=60 periods
Investor yield = YTM = Rate = 7.75%. So Semiannaul yield = 7.75%/2

So HY Coupon payment = PMT(Rate,nper,PV,FV) = PMT(7.75%/2,60,-1107.20,1000)
ie HY PMT = $43.38
So Annual Payment = 2*$43.38 = $86.75
So COupon Rate = PMT/FV = 86.75/1000 = 8.675%

answered by: HApida